# Decision Machines: Enhanced Decision Trees

### v6

@@ Abstract
This paper presents Decision Machines (DMs), an innovative evolution of traditional binary decision trees, which leverages matrix computations to significantly enhance both computational efficiency and interpretability. By explicitly mapping the dependencies between predictions and binary tests within a vector space, DMs offer a streamlined approach to navigating decision paths. We integrate decision trees with kernel methods, ensemble methods and attention mechanisms. The integration of these elements not only bolsters the hierarchical structure of decision trees but also aligns with the computational efficiency of matrix computations. Our work bridges the gap between traditional machine learning algorithms and modern deep learning techniques, providing a novel foundation for further research and application in the field of machine learning.
@@ 1 Introduction
Decision trees have long been a cornerstone in the field of machine learning, revered for their simplicity and interpretability. However, their traditional induction methods, which are inherently greedy and data-driven, often lead to a representation that is less than optimal for computational efficiency. The training process recursively partitions the training space, dividing it into increasingly smaller subsets until a stopping criterion is met, resulting in a tree structure where leaf nodes predict based on the majority of their ground truth set. This approach, while effective, culminates in an implementation that is rule-based and less amenable to computational execution.
@@
In contrast, other supervised learning methods have moved towards end-to-end training schemes, which offer more flexibility and efficiency. The quest for a more computationally friendly representation of decision trees has led to various innovations, including vectorized implementations and attempts to translate decision trees into neural networks. There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [31, 19, 37, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[38]. The oblique decision trees are induced from the derivatives of neural models [22]. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 35, 2]. This paper proposes Decision Machines (DMs), a paradigm that reframes binary decision trees in matrix computation, thereby improving both efficiency and interpretability.
@@
DMs offer a compact representation of binary decision trees, explicitly drawing dependencies between predictions and binary tests. By constructing an arithmetic procedure, we guide input instances from the root to their respective exit leaves, thereby bridging the gap between the intuitive structure of decision trees and the computational efficiency of matrix computations. This formulation not only enhances the interpretability of decision trees but also paves the way for leveraging deep learning toolkits to expedite their execution.
@@
Furthermore, this paper explores the convergence of decision trees with attention mechanisms, a transformative concept in the realm of deep learning. By establishing a connection between tree-based models and attention mechanisms, we open up new avenues for optimizing decision trees and potentially enhancing their predictive power. Our contributions include a provably correct algorithm for representing binary trees analytically, a novel interpretation of binary decision trees as error-correcting output codes, and evidence supporting the derivation of attention mechanisms from decision trees.
@@ 2.1 Construction
We begin with the description of binary decision trees from [25].
@@
For simplicity, we only consider the numerical variables to test, i.e., the input instance $x$ is a real vector in $\mathbb{R}^{n}$. And Boolean conditions are used to test whether the numerical variables are no greater than some given specific thresholds as shown in (1).
In order to represent the true conditions as $-1$s, we substitute 00 with $-1$ in the sign function when $x=0$. And we define the $\operatorname{sgn}(x)$ for $x\in\mathbb{R}$ in our context by
$\operatorname{sgn}(x)=\begin{cases}-1,&\text{if $x\leq 0$},\\ +1,&\text{if $x>0$}.\end{cases}$ (1)
@@
$h(x)=\operatorname{sgn}(Sx-t),$ (2)
the matrix $S$ is of $\{0,1\}^{(L-1)\times n}$ acting as variable selection to test; the instance $x$ is a real vector in $\mathbb{R}^{n}$; the parameter $t$ is a real vector in $\mathbb{R}^{L-1}$; the integer $L$ is the total number of the leaves.
[…]
@@
The template vector of a leaf node in the decision tree is defined by
@@
If the input takes all the tests, it is supposed to follow the ‘maximum likelihood path’. The key step is to compare the similarities between the result vector $h\in\{1,-1\}^{L-1}$ with the template vectors in $\{1,0,-1\}^{L-1}$.
@@
$LS(k,h)=\frac{k\cdot h}{\|k\|_{2}^{2}},$ (3)
where the positive integer $\|k\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $k\in\{1,0,-1\}^{L-1}$. Our similarity(3) is asymmetric and bounded by $(-1,1]$.
@@
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is equal to $1$.
  • If instance $x$ can reach the leaf node $i$ with template vector $k_{i}$, then the similarity $LS(k_{i},h(x))$ is the maximum of all possible paths.
For a binary decision tree, we define its template matrix denoted $B$ in $\{1,0,-1\}^{L\times(L-1)}$ by all the template vectors as its rows. We compute the similarity function (3) on a set of normalized template vectors simultaneously, packed together into a matrix $\bar{B}$ as shown below
$\begin{split}\bar{B}\,h(x))&=(\frac{B_{1}h(x)}{\|B_{1}\|_{2}^{2}},\cdots,\frac{B_{L}h(x)}{\|B_{L}\|_{2}^{2}})^{T}\\ &=(LS(B_{1},h),\cdots,LS(B_{L},h))^{T}\end{split}$ (4)
where the template vector $B_{i}\in\{1,0,-1\}^{L-1}$ is the $i$th row of matrix $B$; the positive integer $\|B_{i}\|_{2}^{2}$ is the squared $\ell_{2}$ norm of the vector $B_{i}$ for $i=1,2,\cdots,L$; the operator $LS(\cdot,\cdot)$ is the logical similarity defined in (3).
@@
$\begin{split}h&=\operatorname{sgn}(Sx-t)\in\{-1,1\}^{L-1},\\ p&=\bar{B}h\in\mathbb{R}^{L},\\ v[i]&=v[\arg\max{(p)}],\end{split}$ (5)
where the function $\arg\max(p)$ returns the index of maximums in the vector $p$; the index $i$ is equal to $\arg\max(p)$; the output $v[i]$ is the value of the $i$-th leaf node for $i=1,2,\cdots,L$. We call the new formulation of binary decision trees logical decision machines.
@@
Suppose the input variable $x$ has 4 numerical features and the trained decision tree is as shown in the figure 1, we will take a specific example to show how logical decision machine(5) works.
The threshold vector is $t=(1,4,3,2,5)^{T}$. The value vector may be categorical or numerical, such as $v=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix and template matrix of the tree 1 are
$S=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix},\quad B=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
Suppose the input instance $x$ is $(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure of logical decision tree is as shown below
$Sx-t=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad h=\operatorname{sgn}(Sx-t)=\begin{pmatrix}+1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},\quad p=\bar{B}h=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
$i=\arg\max(p)=5,v[i]=v_{5}.$
In summary, we obtain $T(x)=v[\arg\max(\bar{B}\,\operatorname{sgn}(Sx-t))]=v_{5}$.
@@ 2.2 Provable guarantee
The template matrix $B$ in (5) is related with the structure of decision tree.
@@
Each leaf node is attached to a row vector in the template matrix $B$. According to definition of template vector, each internal node is attached to a column vector in the template matrix $B$, where its element is attached to a leaf node. The root node will test all samples so that its corresponding results are not zero in all template vectors. In another word, there is a column vector without any 0 in $B$. For example, the first column of $B$ of the decision tree shown in 1 is associated with the test of the root, which means the test is true for the first four leaves and false for the last two ones. Thus, it implies that the first four leaves belongs to left subtree of the root node and the last two ones belong to the right subtree of the root node. Every test will produce two results, so each column must have at least one $1$ and $-1$. In another word, each test must be involved with at least two decision paths.
@@
The template matrix $B$ is of full column-rank.
@@
It is equivalent to prove that the column vectors of $B$ are linearly independent.
For simplicity, we call the column vectors of $B$ in the right subtree as right vectors and the column vectors of $B$ in the left subtree as left vectors.
@@
We want to prove that $LS(b_{i},h(x))=1$ if and only if the instance $x$ can reach the leaf $i$ where the vector $b_{i}$ is the $i$th template vector of a decision tree for $i\in\{1,\cdots,L\}$ and $h(x)$ is the result vector for an input instance $x$,
If the sample $x$ reach the leaf node $i$, it is easy to verify that $LS(b_{i},h(x))=\max_{i\in\{1,2,\cdots,L\}}LS(b_{i},h(x))=1$.
If $LS(b_{i},h(x))=1$, we can obtain $b_{i},h(x)-b_{i}\right>=0$ by the definition (3) so $h(x)$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $x$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. And the maxima in (4) is unique because $LS(b_{i},h(x))\neq LS(b_{j},h(x))$ when $i\neq j$. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
@@ 2.3 Consistent interpretation
We can provide some consistent interpretations of the logical decision machines with other machine learning models.
@@
$\ell_{j}.val=\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}}\frac{y_{i}\delta(\|\vec{t},\vec{t}_{i}\|_{\mathrm{P}})}{N_{j}}$ (6)
where the function $\delta(x)$ is equal to $1$ if $x=0$ and 00 otherwise; $\vec{t}$ and $\vec{t}_{i}$ are the test vectors of new instance $\mathbf{x}$ and the $i$-th input instance respectively; the number $N_{j}$ is the total sum of samples falling in the $j$th leaf node; the notation $\|\vec{t},\vec{s}\|_{\mathrm{P}}$ is a distance defined as $\|\mathrm{P}(\vec{t}-\vec{s})\|_{2}$.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$. The encoding stage of ECOC is described in the algorithms(1).
@@
  • test example $\mathbf{x}$.
  • the coding matrix $M\in\{-1,0,+1\}^{k\times\ell}$ for class $1,2,\cdots,k$.
  • the pre-trained binary classifiers $f=\{f_{s}\mid s=1,2,\cdots,\ell\}$ for each column of $M$.
@@
4:     Compute the binary predictions of test example $\mathbf{x}$: $p=(f_{1}(x),\dots,f_{\ell}(x))$
5:     for  row $r\in M$ do $\triangleright$ iterate over the rows
6:         $d_{r}\leftarrow\mathrm{dis}(r,p)$
@@
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $B$ corresponds to the error-correcting code matrix $M$. All internal nodes are associated with binary test, which are parameterized by $S$ and $t$. And each binary test corresponds to one dichotomizer. In fact, there is a ternary variant of the error-correcting output codes such as [1]. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $B$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree. Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaf nodes in decision trees.

### v7

# Decision Machines: Congruent Decision Trees

@@ Abstract
The decision tree recursively partitions the input space into regions and derives axis-aligned decision boundaries from data. Despite its simplicity and interpretability, decision trees lack parameterized representation, which makes it prone to overfitting and difficult to find the optimal structure. We propose Decision Machines, which embed Boolean tests into a binary vector space and represent the tree structure as a matrices, enabling an interleaved traversal of decision trees through matrix computation. Furthermore, we explore the congruence of decision trees and attention mechanisms, opening new avenues for optimizing decision trees and potentially enhancing their predictive power.
@@ 1 Introduction
Decision trees have long been a cornerstone in the field of machine learning, revered for their simplicity and interpretability. However, their common induction methods, which are inherently greedy and data-driven, often lead to a representation that is less than optimal for computational efficiency. The training process recursively partitions the training space, dividing it into increasingly smaller subsets until a stopping criterion is met, resulting in a tree structure where leaf nodes predict based on the majority of their ground truth set. This approach, while effective, culminates in an implementation that is rule-based and less amenable to machine execution.
@@
In contrast, other supervised learning methods have moved towards end-to-end paradigms, which offer more flexibility and efficiency. The quest for a more computationally friendly representation of decision trees has led to various innovations, including vectorized implementations and attempts to translate decision trees into neural networks. There are some end-to-end training schemes instead of the greedy training methods for decision trees such as [34, 19, 40, 11, 16]. And the implementation of decision trees can be vectorized such as [25, 6, 23, 26] and RapidScorer[41]. The oblique decision trees are induced from the derivatives of neural models [22]. In the section 2 of [9], there is an equivalent model of recursive partitioning in the form of an expansion in a set of basis functions (as products of different step functions) by replacing the geometrical concepts of regions and splitting with the arithmetic notions of adding and multiplying. And it is to transform each tree into a polynomial form in [20]. There are also some schemes to translate the decision trees into neural networks such as [14, 38, 2]. This paper proposes Decision Machines (DMs), a paradigm that reframes binary decision trees in matrix computation, thereby improving both efficiency and interpretability.
@@
DMs offer a compact representation of binary decision trees, explicitly drawing dependencies between predictions and Boolean tests. The binary decision tree is decomposed into two parts: the tree structure representation and the Boolean tests embedding. The binary tree structure is represented by a character matrix and the evaluation of Boolean tests are embedded into a binary vector space. DMs navigate the input insatnce from the root to its destined exit leaf based on the mutiplication of the tree matrix and the embedding vector. This formulation not only enhances the interpretability of decision trees but also paves the way for leveraging deep learning toolkits to expedite their execution.
@@
Furthermore, this paper explores the congruence of decision trees and attention mechanisms, a revolutionary concept in the realm of deep learning, in order to open up new avenues for optimizing decision trees and potentially enhance their predictive power. Our contributions include a provable algorithm for representing binary trees analytically, a novel interpretation of binary decision trees as error-correcting output codes, and some evidence supporting the derivation of attention mechanisms from decision trees.
@@ 2.1 Construction
We adopt the description of binary decision trees from [25].
@@
For simplicity, we only consider the numerical variables to test, i.e., the input instance $\mathbf{x}$ is a real vector in $\mathbb{R}^{n}$. And Boolean conditions are used to test whether the numerical variables are no greater than some given specific thresholds.
In order to represent the true conditions as $-1$s, we substitute 00 with $-1$ in the sign function when 00. And we define the $\operatorname{sgn}(z)$ for $z\in\mathbb{R}$ in our context by
$\operatorname{sgn}(z)=\begin{cases}-1,&\text{if $z\leq 0$},\\ +1,&\text{if $z>0$}.\end{cases}$ (1)
@@
$h(\mathbf{x})=\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),$ (2)
the matrix $\mathrm{S}$ is of $\{0,1\}^{(L-1)\times n}$ acting as variable selection to test; the instance $\mathbf{x}$ is a real vector in $\mathbb{R}^{n}$; the parameter $\mathbf{t}$ is a real vector in $\mathbb{R}^{L-1}$; the integer $L$ is the total number of the leaves.
$\mathbf{x}[1]\leq 1?$ $\mathbf{x}[2]\leq 4?$$\mathbf{x}[3]\leq 3?$$\mathbf{x}[2]\leq 2?$$v_{4}$$v_{5}$$v_{6}$$v_{1}$$\mathbf{x}[4]\leq 5?$$v_{2}$$v_{3}$rueFalseTrueFalseTrueFalseTrueFalseTrueFalse
@@
The template vector of a leaf node is defined by the following steps:
@@
If the input takes all the tests, it is supposed to follow the ‘maximum likelihood path’. The key step is to compare the similarities between the result vector $\mathbf{h}\in\{1,-1\}^{L-1}$ with the template vectors in $\{1,0,-1\}^{L-1}$.
@@
$LS(\mathbf{k},\mathbf{h})=\frac{\mathbf{k}\cdot\mathbf{h}}{\|\mathbf{k}\|_{2}^{2}},$ (3)
where the positive integer $\|k\|_{2}^{2}$ is the squared $\ell_{2}$ norm of template vector $\mathbf{k}\in\{1,0,-1\}^{L-1}$. Our similarity(3) is asymmetric and bounded by $(-1,1]$.
@@
  • If instance $\mathbf{x}$ can reach the leaf node $i$ with template vector $\mathbf{k}$, then the similarity $LS(\mathbf{k},h(\mathbf{x}))$ is equal to $1$.
  • If instance $\mathbf{x}$ can reach the leaf node $i$ with template vector $\mathbf{k}$, then the similarity $LS(\mathbf{k},h(\mathbf{x}))$ is the maximum of all possible paths.
For a binary decision tree, the template matrix $\mathrm{B}$ consists of all the template vectors as its rows. We compute the similarity function (3) on a set of normalized template vectors simultaneously, packed together into a matrix $\bar{\mathrm{B}}$ as shown below
$\begin{split}\bar{\mathrm{B}}\,h(\mathbf{x})&=(\frac{B_{1}h(\mathbf{x})}{\|B_{1}\|_{2}^{2}},\cdots,\frac{B_{L}h(\mathbf{x})}{\|B_{L}\|_{2}^{2}})^{T}\\ &=(LS(B_{1},h),\cdots,LS(B_{L},h))^{T}\end{split}$ (4)
where the template vector $B_{i}\in\{1,0,-1\}^{L-1}$ is the $i$th row of matrix $\mathrm{B}$; the positive integer $\|B_{i}\|_{2}^{2}$ is the squared $\ell_{2}$ norm of the vector $B_{i}$ for $i=1,2,\cdots,L$; the operator $LS(\cdot,\cdot)$ is the logical similarity defined in (3).
@@
$\begin{split}\mathbf{h}&=\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t}),\\ \mathbf{p}&=\bar{\mathrm{B}}\mathbf{h}\in\mathbb{R}^{L},\\ v[i]&=v[\arg\max{(\mathbf{p})}],\end{split}$ (5)
where the function $\arg\max(\mathbf{p})$ returns the index of maximums of $\mathbf{p}$; the index $i$ is equal to $\arg\max(\mathbf{p})$; the output $v[i]$ is the value of the $i$-th leaf node for $i=1,2,\cdots,L$. We call the new formulation of binary decision trees logical decision machines.
@@
Suppose the input variable $\mathbf{x}$ has 4 numerical features and the trained decision tree is as shown in the figure 1, we will take a specific example to show how logical decision machine(5) works.
The threshold vector is $\mathbf{t}=(1,4,3,2,5)^{T}$. The value vector may be categorical or numerical, such as $\mathbf{v}=(v_{1},v_{2},v_{3},v_{4},v_{5},v_{6})^{T}$. The selection matrix and template matrix of the tree 1 are
$\mathrm{S}=\begin{pmatrix}1&0&0&0\\ 0&1&0&0\\ 0&0&1&0\\ 0&1&0&0\\ 0&0&0&1\end{pmatrix},\quad\mathrm{B}=\begin{pmatrix}-1&-1&0&-1&0\\ -1&-1&0&+1&-1\\ -1&-1&0&+1&+1\\ -1&+1&0&0&0\\ +1&0&-1&0&0\\ +1&0&+1&0&0\\ \end{pmatrix}.$
Suppose the input instance $\mathbf{x}$ is $(2,1,2,2)^{T}$, its prediction is $v_{5}$ according to the figure 1. The computation procedure of logical decision tree is as shown below
$\mathrm{S}\mathbf{x}-\mathbf{t}=\begin{pmatrix}1\\ -3\\ -1\\ -1\\ -3\end{pmatrix},\quad\mathbf{h}=\operatorname{sgn}(\mathrm{S}\mathbf{x}-\mathbf{t})=\begin{pmatrix}1\\ -1\\ -1\\ -1\\ -1\end{pmatrix},\quad\mathbf{p}=\bar{\mathrm{B}}\mathbf{h}=\begin{pmatrix}\frac{1}{3}\\ 0\\ -\frac{1}{2}\\ 0\\ 1\\ 0\end{pmatrix},$
$i=\arg\max(\mathbf{p})=5,\mathbf{v}[i]=v_{5}.$
In summary, we obtain $T(\mathbf{x})=v_{5}$.
@@ 2.2 Provable guarantee
The template matrix $\mathrm{B}$ in (5) is related with the structure of decision tree.
@@
Each leaf node is attached to a row vector in the template matrix $\mathrm{B}$. According to definition of template vector, each internal node is attached to a column vector in the template matrix $\mathrm{B}$, where its element is attached to a leaf node. The root node will test all samples so that its corresponding results are not zero in all template vectors. In another word, there is a column vector without any 0 in $\mathrm{B}$. For example, the first column of $\mathrm{B}$ of the decision tree shown in 1 is associated with the test of the root, which means the test is true for the first four leaves and false for the last two ones. Thus, it implies that the first four leaves belongs to left subtree of the root node and the last two ones belong to the right subtree of the root node. Every test will produce two results, so each column must have at least one $1$ and $-1$. In another word, each test must be involved with at least two decision paths.
@@
The template matrix $\mathrm{B}$ is of full column-rank.
@@
It is equivalent to prove that the column vectors of $\mathrm{B}$ are linearly independent.
For simplicity, we call the column vectors of $\mathrm{B}$ in the right subtree as right vectors and the column vectors of $\mathrm{B}$ in the left subtree as left vectors.
@@
We want to prove that $LS(b_{i},h(\mathbf{x}))=1$ if and only if the instance $\mathbf{x}$ can reach the leaf $i$ where the vector $b_{i}$ is the $i$th template vector of a decision tree for $i\in\{1,\cdots,L\}$ and $h(\mathbf{x})$ is the result vector for an input instance $\mathbf{x}$,
If the sample $\mathbf{x}$ reach the leaf node $i$, it is easy to verify that $LS(b_{i},h(\mathbf{x}))=\max_{i\in\{1,2,\cdots,L\}}LS(b_{i},h(\mathbf{x}))=1$.
If $LS(b_{i},h(\mathbf{x}))=1$, we can obtain $b_{i},h(\mathbf{x})-b_{i}\right>=0$ by the definition (3) so $h(\mathbf{x})$ is equal to $b_{i}$ for every non-zero element in $b_{i}$, which means the instance $\mathbf{x}$ can satisfy the conditions of the $i$th leaf’s ancestor nodes. Becasue the exit leaf node is unique for every sample, so the the maxima in (4) is unique. Thus, the maxima in (4) is sufficient to identify the exit leaf. And we complete the proof.
@@ 2.3 Consistent interpretation
The logical decision machines are consistent with some machine learning models.
@@
$\ell_{j}.val=\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{D}}\frac{y_{i}\delta(\|\mathbf{h},\mathbf{h}_{i}\|_{\mathrm{B}})}{N_{j}}$ (6)
where the function $\delta(z)$ is equal to $1$ if $z=0$ and 00 otherwise; $\mathbf{h}$ and $\mathbf{h}_{i}$ are the hidden vectors of new instance $\mathbf{x}$ and $\mathbf{x}_{i}$ respectively; the number $N_{j}$ is the total sum of samples falling in the $j$th leaf node; the matrix $\mathrm{B}$ is the template matrix; the notation $\|\mathbf{h},\mathbf{k}\|_{\mathrm{B}}$ is a distance defined as $\|\mathrm{B}(\mathbf{h}-\mathbf{k})\|_{2}$.
The algorithm 5 is also similar to the error correcting output codes (ECOC) such as [7, 18, 1], which is a powerful tool to deal with multi-class categorization problems. The basic idea of ternary ECOC is to associate each class $r$ with a row of a coding matrix $\mathrm{M}\in\{-1,0,+1\}^{k\times\ell}$. The encoding stage of ECOC is described in the algorithms(1).
@@
  • the test example $\mathbf{x}$.
  • the coding matrix $\mathrm{M}\in\{-1,0,+1\}^{k\times\ell}$ for class $1,2,\cdots,k$.
  • the pre-trained binary classifiers $f=\{f_{s}\mid s=1,2,\cdots,\ell\}$ for each column of $\mathrm{M}$.
@@
4:     Compute the binary predictions of $\mathbf{x}$: $\mathbf{p}=(f_{1}(\mathbf{x}),\dots,f_{\ell}(\mathbf{x}))$
5:     for  row $\mathbf{r}\in\mathrm{M}$ do $\triangleright$ iterate over the rows
6:         $d_{r}\leftarrow\mathrm{dis}(\mathbf{r},\mathbf{p})$
@@
The bit-vector corresponds to the error-correcting code $h$ and the bit-vector matrix $\mathrm{B}$ corresponds to the error-correcting code matrix $\mathrm{M}$. All internal nodes are associated with binary test, which are parameterized by $\mathrm{S}$ and $\mathbf{t}$. The main difference between 1 and (5) is that the row in the code matrix $M$ corresponds to a specific class while the row in the template matrix $\mathrm{B}$ corresponds to a specific leaf. In one hand, we can consider our method as another coding strategy of ternary error-correcting output codes. In another hand, the coding strategy such as [8] will help to find new induction methods of decision tree. Their main difference is that each class is attached with a single code word in ECOC while each class may be associated with multiple leaf nodes in decision trees.